OpenStudy (katy):

So, this math problem has be solve for the first, second, and third derivatives of an equation and then asks me to "guess a formula for d''y/dx''" I get derivatives, but I don't get what the question is asking. Wouldn't the formula just be what I found for y''? Help, please! Here's the given equation and the three derivatives I found, if that helps any: y=1/(1-x); y'=1/(1-x)^2 y''=2/(1-x)^3 y'''=6/(1-x)^4

OpenStudy (anonymous):

you want a hint? denominator is pretty clear, question is, what is the numerator. derivative numerator 1 1 2 2 3 6 4 24 5 120 if the right hand is not familiar, think of where the numbers come form 6 comes from \(3\times 2\) 24 comes from \(4\times 6\) i.e. \(4\times 3\times 2\) etc

OpenStudy (eyust707):

\[1 = 1 \] \[1*2 = 2\] \[1*2*3 = 6\] \[1 * 2 * 3 * 4 = 24\] \[1 * 2 * 3 * 4 * 5 = 120\]

OpenStudy (anonymous):

if i read the question correctly you are supposed to recognize the numerators as a factorial

OpenStudy (katy):

So I see the relationship now, but how is that going to become formula different from I already found for y''?

OpenStudy (katy):

Theses answers really aren't helping me yet.